OpenStudy (anonymous):

How many subsets of at least four elements does a set of seven elements have?

OpenStudy (anonymous):

7C4 +7C5 + 7C6 +7C7

OpenStudy (anonymous):

use calculator

OpenStudy (anonymous):

for sure.

OpenStudy (anonymous):

thank you again elecengineer

OpenStudy (anonymous):

or not. 1 with 7 elements 7 with 6 elements \[\dbinom{7}{5}=\frac{7\times 6}{2}=21\] with 5 elements and \[\dbinom{7}{4}=\frac{7\times 6\times 5}{3\times 2}=35\] for 4 elements

OpenStudy (anonymous):

\[1+7+21+35=64\]

OpenStudy (anonymous):

thank you satellite!